GDP 1000 Government Expenditure 120 Investment Expenditure 160 Consumption Expenditure 700 Imports 80 1) Use the data in the table above to answer the following questions. a) [3 points] Based on the data provided above, what is the value of National Savings? b) [3 points] Is this country a lender or a borrower? c) [3 points] What is the value of exports?
Answer #1

a) Ans: 180

Explanation:

National Savings = GDP - C - G = 1000 - 700 - 120 = 180

b) Ans: Lender

Explanation:

Domestic Savings = National Savings = 180

Domestic Investment = 160

Trade Surplus = Domestic Savings - Domestic Investment = 180 - 160 = 20

So, the country is a lender.

c) Ans: 100

Explanation:

GDP = C + I + G + X - M

1000 = 700 + 160 + 120 + X - 80

1000 = 900 + X

X = 1000 - 900 = 100

Thus, Export is 100.